### v2

# MoM: Linear Sequence Modeling with Mixture-of-Memories

Abstract
Linear sequence modeling methods, such as linear attention, state space modeling, and linear RNNs, offer significant efficiency improvements by reducing the complexity of training and inference. However, these methods typically compress the entire input sequence into a single fixed-size memory state, which leads to suboptimal performance on recall-intensive downstream tasks. Drawing inspiration from neuroscience, particularly the brain’s ability to maintain robust long-term memory while mitigating "memory interference", we introduce a novel architecture called Mixture-of-Memories (MoM). MoM utilizes multiple independent memory states, with a router network directing input tokens to specific memory states. This approach greatly enhances the overall memory capacity while minimizing memory interference. As a result, MoM performs exceptionally well on recall-intensive tasks, surpassing existing linear sequence modeling techniques. Despite incorporating multiple memory states, the computation of each memory state remains linear in complexity, allowing MoM to retain the linear-complexity advantage during training, while constant-complexity during inference. Our experimental results show that MoM significantly outperforms current linear sequence models on downstream language tasks, particularly recall-intensive tasks, and even achieves performance comparable to Transformer models.
The code is released at https://github.com/OpenSparseLLMs/MoM and is also released as a part of https://github.com/OpenSparseLLMs/Linear-MoE.
MoM: Linear Sequence Modeling with Mixture-of-Memories
Jusen Du1,2, Weigao Sun1🖂, Disen Lan1,3, Jiaxi Hu4, Yu Cheng5🖂 1Shanghai AI Laboratory, 2Nanjing University, 3South China University of Technology, 4The Hong Kong University of Science and Technology (Guangzhou), 5The Chinese University of Hong Kong
"The enemy of memory isn’t time; it’s other memories."
— David Eagleman
† interns at Shanghai AI Laboratory; 🖂  corresponding authors (sunweigao@outlook.com, chengyu@cse.cuhk.edu.hk). Weigao Sun is the project lead.
1 Introduction
Attention mechanisms have made significant contributions to the field of artificial intelligence, advancing various modalities such as language, vision, audio, video, graphs, and even time series Achiam et al. (2023); Team (2023). The Transformer Vaswani (2017), known for its ability to capture long-range dependencies, has become a foundational architecture in this space. However, traditional Transformers encounter computational challenges due to their quadratic time complexity, $O(n^{2})$, with respect to sequence length $n$, making it difficult to scale to long sequences. To overcome this limitation, several linear sequence modeling methods have been proposed, including linear attention Katharopoulos et al. (2020); Qin et al. (2023a); Li et al. (2025), state space modeling Gu and Dao (2024); Dao and Gu (2024), and linear RNNs Peng et al. (2024); Qin et al. (2024d), which offer $O(n)$ training complexity and $O(1)$ inference complexity. These approaches often reduce the input sequence to a fixed-size hidden space, collapsing the information into a single "memory state". While these methods enhance efficiency, they face two main challenges: limited memory capacity and memory interference. When new information overwrites the single fixed-size memory state, previously stored representations may degrade, which negatively impacts its long-term memory performance on recall-intensive tasks.
We argue that the strong performance of Transformer models on recall-intensive tasks arises from their ability to avoid "memory interference" by maintaining independent key-value caches for each token, thus offering virtually unlimited memory capacity. In contrast, linear sequence modeling relies on extreme compression, consolidating all the input information into a single fixed-size memory state Katharopoulos et al. (2020); Dao and Gu (2024). This approach results in limited memory capacity and inherently leads to memory interference issues.
Interestingly, the human brain has developed mechanisms that enable large memory capacity while reducing memory interference. Neuroscience studies show that in the hippocampus, theta oscillations (4$\sim$8 Hz) and gamma oscillations (30$\sim$100 Hz) work together to support a neural coding mechanism for multi-item memory Buzsáki (2002); Lisman and Jensen (2013). Specifically, each theta cycle is subdivided into multiple gamma subcycles, and within each gamma subcycle, a distinct group of neurons is activated following the "E%-max" mechanism de Almeida et al. (2009). This sequential activation temporally separates different memory items, thus preventing interference.
Inspired by these biological insights, we propose a new architecture called Mixture-of-Memories (MoM), which aims to strike a balance between the explicit token representations in Transformers and the extreme compression found in earlier linear sequence modeling methods. MoM employs multiple independent memory states, with a router network that directs input tokens to specific memory states. The input sequence is divided into a predefined number of subsequences (phase-specific neural assemblies), which are processed in parallel and fed into the corresponding memory projections (dentate microcircuits) to generate key-value pairs. As the linear sequence modeling layer processes each subsequence using an RNN-like update mechanism, it produces multiple memory states that capture different aspects of the input sequence. The final output is computed as a weighted sum of these memories, which we refer to as the mixture-of-memories. This approach expands memory capacity and eliminates memory interference, enabling MoM to significantly outperform existing linear sequence models that rely on a single fixed-size memory state.
Our contributions can be summarized as follows:
  • We present MoM, a neuroscience-inspired architecture that incorporates multiple independent memory states, significantly enhancing memory capacity and eliminating memory interference, while retaining the efficiency benefits of linear-time training and constant-memory inference.
  • Distinct with existing gating mechanisms, MoM is a new paradigm to reduce memory interference by separating the memory states. The overall design of MoM is compatible with existing linear sequence modeling methods, making it a straightforward and effective approach to boost task performance.
  • Through empirical evaluation, we show that MoM outperforms state-of-the-art linear sequence modeling techniques across a variety of language tasks, particularly on recall-intensive tasks. MoM even achieves performance on par with Transformer models, a feat that current linear sequence modeling methods struggle to match.
2 Preliminary
For notations in this work, we use bold lower-case letters for row vectors (e.g., $\bm{q}_{t},\bm{k}_{t}$), bold upper-case letters for matrices (e.g., $\bm{Q},\bm{K}$) and the identical letters represent a row in the matrix, e.g., $\bm{q}_{t}$ is the $t$-th row of $\bm{Q}$.
Linear Attention
To reduce the time complexity of Transformer attention, various optimization techniques have been proposed. Linear Transformers Katharopoulos et al. (2020) replace the softmax attention mechanism with dot-product of feature maps $\phi(\cdot)$:
$\bm{o}_{t}=\frac{\sum_{i=1}^{n}\phi(\bm{q}_{t})\phi(\bm{k}_{i})^{T}\bm{v}_{i}}{\sum_{i=1}^{n}\phi(\bm{q}_{t})\phi(\bm{k}_{i})^{T}},$ (1)
where $\bm{q}_{t},\bm{k}_{t},\bm{v}_{t}\in\mathbb{R}^{d}$. While the presence of the denominator may lead to numerical instability Qin et al. (2024b) and the feature map can utilize an identity function, which we omit for simplicity. In perspective of memory, the formulation can also be written in a recurrent format:
$\bm{M}_{t}=\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t},\quad\bm{o}_{t}=\bm{q}_{t}\bm{M}_{t},$ (2)
This indicates that linear attention can function as a linear recurrent layer with a matrix-valued hidden state $\bm{M}$ which we refer to as memory sate and the output is generated by querying the memory state $\bm{M}$. This represents the ultimate compression of sequence information, condensing the entire sequence into a single memory state.
Building on the foundational concepts of linear attention and memory perspective, some recent advancements have focused on optimizing memory structure, including gated updates Yang et al. (2023); Qin et al. (2024e, d) and memory capacity expansion Peng et al. (2024); Qin et al. (2024d).
Figure 1: MoM Architecture. Each input token selectively activates and updates $K$ memory states, leaving non-activated memory states unchanged to avoid interference from current input. Additionally, we introduce a continuously activated shared memory. This figure presents the basic memory update mechanism; other mechanisms involving gating or more complex updates follow a similar approach.
3 Method
3.1 Motivation
Linear sequence models compress the entire sequence data into a fixed-size memory state. Despite numerous efforts to minimize information loss—such as introducing gating mechanisms and employing more precise control over memory modifications Orvieto et al. (2023); De et al. (2024); Beck et al. (2024); Yang et al. (2023); Zhang et al. (2024)—some degradation in this compression process is inevitable. Expanding the memory capacity has been shown to mitigate this issue to some extent, with studies indicating that increasing memory capacity can enhance model performance Qin et al. (2024d); Peng et al. (2024).
However, previous approaches that simply increased the size of the RNN state, essentially expanding a single memory state, struggled to capture the full spectrum of information within an entire sequence. We propose that this difficulty arises because sequence information is often multifaceted, and a single, expanded memory may not be capable of simultaneously capturing multiple aspects of the data. Inputs that introduce new or orthogonal information may interfere with existing memory content when using a shared memory. Rather than discarding these inputs through gating mechanisms or overwriting the existing memory state, it may be more effective to consider alternative strategies that allow for the preservation of diverse information without interference.
3.2 MoM: Mixture-of-Memories
To address the challenge outlined above, we propose a novel approach inspired by biological mechanisms for encoding multi-item memory such as theta-gamma oscillations Lisman and Jensen (2013), and concepts from Mixture-of-Experts (MoE) Shazeer et al. (2017), where different experts handle specific tokens. In this approach, we leverage multiple memory states, each of which is selectively updated by different inputs. This increases the memory capacity and enables the model to retain diverse pieces of information by storing various types of inputs in separate memory states.
In our framework, the memory states function similarly to the experts in MoE. However, instead of relying on completely separate networks, these modules are individual RNN states embedded within a linear recurrent mechanism. This design allows for the isolation of memory updates while concurrently managing distinct types of information. It is important to note that MoM fundamentally differs from traditional MoE, as we will discuss in Appendix B. Figure 1 provides an overview of the MoM architecture. Below, we introduce the structure of the MoM layer and explain how this biologically-inspired architecture is implemented in the context of linear sequence modeling.
3.2.1 Router
We use a router to assign inputs to different memory states. Utilizing the top-$k$ concept, each token is routed to the top-$k$ memories based on its importance scores. Specifically, we use a simple linear layer to generate these scores for each input token. After applying a softmax function, we select the top-$k$ scores and normalize them.
$\mathbf{scores}_{t}=\mathrm{TopK}(\mathrm{softmax}(\bm{x}_{t}\bm{W}_{g}))\in\mathbb{R}^{k},$ (3)
$\bm{g}_{t}=\frac{\mathbf{scores}_{t}}{\sum\mathbf{scores}_{t}}\in\mathbb{R}^{k},$ (4)
where $\bm{x}_{t}\in\mathbb{R}^{d}$, $k$ is the top-$k$ number, $\bm{W}_{g}\in\mathbb{R}^{d\times M}$ is learnable weight, $\bm{g}_{t}$ is the normalized importance scores of the input $\bm{x}_{t}$.
3.2.2 Linear Recurrent Memory Module
After the router network, the input $\bm{x}_{t}$ is directed to top-$k$ linear recurrent modules, meaning that the top-$k$ memories are activated while the others remain inactive.
Each Memory. For each activated memory, indexed by $m$, we perform the following operation:
  1. 1.
    Key and Value Projections: We project the input $\bm{x}_{t}$ to $\bm{k}^{m}_{t}$ and $\bm{v}^{m}_{t}$ using $\bm{W}^{m}_{k}$ and $\bm{W}^{m}_{v}$:
    $\bm{k}^{m}_{t}=\bm{x}_{t}\bm{W}^{m}_{k},\bm{v}^{m}_{t}=\bm{x}_{t}\bm{W}^{m}_{v}\in\mathbb{R}^{d},$ (5)
    where $\bm{W}^{m}_{k}$, $\bm{W}^{m}_{v}$ are learnable projection weights for $kv$ of the $m$-th memory module.
  2. 2.
    Memory Update: We update the activated memory state using $\bm{k}^{m}_{t}$, $\bm{v}^{m}_{t}$:
    $\bm{M}^{m}_{t}=\bm{M}^{m}_{t-1}+(\bm{k}^{m}_{t})^{T}\bm{v}^{m}_{t}\in\mathbb{R}^{d\times d},$ (6)
    The equation above represents the simplest form of memory update for clarity. Our approach is flexible and does not rely on a specific memory update mechanism. To enhance performance, we can incorporate mechanisms such as forget gates Sun et al. (2023):
    $\bm{M}^{m}_{t}=\gamma\bm{M}^{m}_{t-1}+(\bm{k}^{m}_{t})^{T}\bm{v}^{m}_{t}\in\mathbb{R}^{d\times d},$ (7)
    where $\gamma$ is a constant forget gate.
    More generally, our method can be adapted to incorporate various memory update methods proposed in previous work. Detailed descriptions of these methods are provided in Table 1.
Method Memory Update Rule
Linear Attn $\bm{M}_{t}=\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$
Lightning $\bm{M}_{t}=\gamma\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$
RetNet $\bm{M}_{t}=\gamma\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$
GLA $\bm{M}_{t}=(\bm{a}_{t}^{T}\bm{1})\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$
DeltaNet $\bm{M}_{t}=(\bm{I}-\bm{k}_{t}^{T}\bm{k}_{t})\bm{M}_{t-1}+b_{t}\bm{k}_{t}^{T}\bm{v}_{t}$
G-DeltaNet $\bm{M}_{t}=a_{t}(\bm{I}-\bm{k}_{t}^{T}\bm{k}_{t})\bm{M}_{t-1}+b_{t}\bm{k}_{t}^{T}\bm{v}_{t}$
TTT $\bm{M}_{t}=\bm{M}_{t-1}+b_{t}\nabla l(\bm{M}_{t-1};\bm{k}_{t},\bm{v}_{t})$
Titans* $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+b_{t}\nabla l(\bm{M}_{t-1};\bm{k}_{t},\bm{v}_{t})$
Mamba2 $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+b_{t}\bm{k}_{t}^{T}\bm{v}_{t}$
HGRN2 $\bm{M}_{t}=(\bm{a}_{t}^{T}\bm{1})\bm{M}_{t-1}+(1-\bm{a}_{t})^{T}\bm{v}_{t}$
RWKV6 $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$
RWKV7 $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+b_{t}\nabla l(\bm{M}_{t-1};\bm{k}_{t},\bm{v}_{t})$
    For Titans, we have omitted details like momentum and norm optimization during gradient descent to keep the expression simplified.
Table 1: Memory Update Rules. We demonstrate that several current linear sequence models can be viewed as recurrent models in terms of memory updates, where $a_{t},b_{t}\in(0,1)$ are data-dependent scaler, $\bm{a}_{t}$ is data-dependent vector, and $\gamma$ is a data-independent constant.
Memory Mixing. After updating the activated memory states, we perform a weighted sum of these memory states using the importance scores obtained from Equation(4).
$\tilde{\bm{M}}_{t}=\sum_{m}g_{t}^{(m)}\bm{M}^{m}_{t}\in\mathbb{R}^{d\times d},$ (8)
where $\bm{M}_{m}$ is one activated memory and $g_{t}^{(m)}$ is the importance score of $\bm{M}_{m}$.
We then obtain the output of the MoM by applying query vector $\bm{q}_{t}$ to the mixed memory $\tilde{\bm{M}}_{t}$:
$\bm{o}_{t}=\bm{q}_{t}\tilde{\bm{M}}_{t}\in\mathbb{R}^{d},$ (9)
Finally, the output of the MoM layer is computed by applying an activation function, normalization, and a linear transformation:
$\bm{o}_{t}=\mathrm{RMSNorm}(\mathrm{Swish}(\bm{o}_{t}))\bm{W}_{o}\in\mathbb{R}^{d},$ (10)
Throughout the recurrent process, only a subset of memory states is activated and updated at each time step, while memory states that are not routed remain inactive and unchanged. When the input passes through the key-value projection layer, it generates multiple sets of keys and values that are fed into different memory modules. This design enables the model to maintain multiple memory states, each preserving distinct pieces of information. By aggregating the activated memories into a comprehensive mixed memory by weighted summation, the query can effectively retrieve information from this mixed memory, which results the "attention output" followed by other layers.
Scale Model FDA SWDE SQUAD NQ TriviaQA Drop Avg.
Avg.
(no FDA)
340M Params 15B Tokens L=24, d=1024 Transformer++ 46.14 25.87 33.22 18.94 45.97 20.03 31.70 28.81
RetNet 5.90 9.28 22.41 6.91 40.05 18.59 17.19 19.45
HGRN2 11.53 17.34 24.08 12.67 43.84 17.35 21.14 23.06
GLA 11.26 16.78 27.85 12.77 43.90 17.68 21.71 23.80
GSA 6.36 16.87 21.90 14.60 42.18 16.72 19.77 22.45
Gated DeltaNet 20.53 23.24 28.55 14.98 44.91 16.48 24.78 25.63
MoM 22.98 29.90 29.69 16.60 48.82 20.99 28.16 29.20
1.3B Params 100B Tokens L=24, d=2048 Transformer++ 44.32 32.43 42.59 24.49 58.47 21.56 37.31 35.91
RetNet 13.62 22.59 33.46 15.43 53.79 19.79 26.45 29.01
HGRN2 12.35 23.24 33.19 19.10 55.27 19.65 27.13 30.09
GLA 27.61 30.93 35.04 22.27 56.28 19.45 31.93 32.79
GSA 23.25 32.80 35.57 22.96 57.05 20.65 32.05 33.81
Gated DeltaNet 30.25 27.65 34.06 23.22 58.23 20.36 32.30 32.70
MoM 41.14 34.30 37.08 24.11 58.59 21.03 36.04 35.02
Table 2: Results on Recall-Intensive Tasks. All inputs are truncated to a maximum length of 2K tokens. MoM significantly outperforms all other linear models across both model sizes. In the 1.3B model, MoM even achieves performance very close to that of Transformer models.
3.3 Optimizations for MoM
Shared Memory.  To enhance our model’s ability to capture long-term dependencies, we introduce a shared memory mechanism. This shared memory has access to the entire sequence information, allowing it to effectively store and retrieve long-term information. By integrating shared memory into our model, we ensure that it can leverage the complete historical context, resulting in significant improvements in performance and robustness.
$\bm{k}^{s}_{t}=\bm{x}_{t}\bm{W}^{s}_{k},\bm{v}^{s}_{t}=\bm{x}_{t}\bm{W}^{s}_{v}\in\mathbb{R}^{d}$ (11)
$\bm{M}^{s}_{t}=\bm{M}^{s}_{t-1}+(\bm{k}^{s}_{t})^{T}\bm{v}^{s}_{t}\in\mathbb{R}^{d\times d}$ (12)
$\tilde{\bm{M}}_{t}=\bm{M}^{s}_{t}+\sum_{m}g_{t}^{(m)}\bm{M}^{m}_{t}\in\mathbb{R}^{d\times d}$ (13)
Hardware-efficient Implementation.   Chunk-wise parallel form of linear attention is a computational optimization strategy that divides the input sequence into smaller chunks to enable partial parallel and recursive computation during model training Hua et al. (2022); Yang et al. (2023), which is well supported in open-source frameworks.
In the implementation of MoM, mixing the memories before multiplying by the query is mathematically equivalent to firstly multiplying each memory by the query and then mixing the results. This property allows us to reuse the efficient Triton-based operator implementation of all previous linear sequence modeling methods, facilitate the hardware-efficient training as well as inference of MoM.
Models marked with an asterisk use open-source pretrained weights with identical training configurations.
Scale Model
Wiki.
$\downarrow$
Lamb.
$\downarrow$
ARC-e
$\uparrow$
ARC-c
$\mathrm{acc_{n}}$$\uparrow$
Hella.
$\mathrm{acc_{n}}$$\uparrow$
Lamb.
$\mathrm{acc}$$\uparrow$
PIQA
$\mathrm{acc}$$\uparrow$
Wino.
$\mathrm{acc}$$\uparrow$
Avg.
340M Params 15B Tokens L=24, d=1024 Transformer++ 26.88 76.46 44.91 25.94 34.95 26.90 64.31 51.07 41.35
RetNet 31.07 87.11 44.49 23.04 33.86 23.93 63.49 52.33 40.19
HGRN2 27.90 77.40 45.24 23.63 35.61 24.74 65.45 54.06 41.46
GLA 28.78 79.95 44.53 22.27 34.84 24.94 63.93 51.38 40.32
GSA 28.17 82.50 45.50 24.23 35.00 24.02 64.85 50.43 40.67
Gated DeltaNet 26.47 58.59 46.04 23.55 35.18 27.01 66.05 50.83 41.44
MoM 25.86 55.41 44.65 24.74 36.54 27.93 66.16 51.78 41.97
1.3B Params 100B Tokens L=24, d=2048 Transformer++ 17.61 19.29 55.01 28.07 49.21 40.95 70.08 56.27 49.93
RetNet 18.18 21.97 57.49 26.88 48.09 37.75 69.37 53.28 48.81
HGRN2 17.32 15.65 58.33 28.07 51.93 42.31 71.33 52.01 50.66
GLA 17.61 19.66 55.18 27.56 48.89 40.03 69.86 53.91 49.24
GSA 16.69 16.02 58.33 28.33 50.98 42.03 72.25 53.43 50.89
Gated DeltaNet 17.14 18.80 56.82 27.39 49.77 39.94 71.76 51.78 49.58
MoM 16.64 14.83 55.35 27.99 50.95 43.43 71.27 56.83 50.97
Table 3: Results on Common-Sense Reasoning Tasks. The performance of linear models and Transformer models is comparable; however, MoM consistently achieves the best average performance across all model sizes.
4 Experiments
4.1 Experimental Setups
Models.   In our experiments, we employ the Gated DeltaNet Yang et al. (2024) as the memory update mechanism in MoM. The model is configured with four memory states, two of which are activated at each time step, along with a shared memory.
Baselines.   We evaluate MoM against several linear recurrent models and Transformers, including RetNet Sun et al. (2023), GLA Yang et al. (2023), Gated DeltaNet Yang et al. (2024), and Transformer++ Touvron et al. (2023), which incorporates Rotary Position Embeddings Su et al. (2024) and GLU Shazeer (2020) into the Transformer architecture. To ensure a fair comparison, we train all baseline models from scratch using the exact same number of tokens.
Training.   We follow the training procedure described by Yang et al. (2023), utilizing the SlimPajama dataset Soboleva et al. (2023) sampled with 100B tokens and tokenized using the Mistral tokenizer Jiang et al. (2023). We train models from scratch with parameter sizes of 340M and 1.3B, respectively. For the 340M models, we train on 15B tokens with a batch size of 0.5M tokens. The warmup tokens count is set to 0.25M. We set the hidden ratio of our model to 3 to keep the activated parameter count approximately the same. For the 1.3B models, we train on 100B tokens with a batch size of 2M tokens. We utilized publicly available pretrained weights from Zhang et al. (2024) with exactly same configuration. The warmup tokens count is 1B. We employ AdamW optimizer Loshchilov et al. (2017); Sun et al. (2024b) with learning rate of 3e-4 with cosine learning rate schedule Zhou et al. (2020). The weight decay is set to 0.01 and gradient clipping is 1.0.
Parameter Explanation. The parameter counts differ across models due to the varying modules they contain. For reference, the "340M" model refers to a configuration with 24 layers, each having a hidden size of 1024. Similarly, "1.3B" refers to 24 layers with a hidden size of 2048. Actual parameter variations may occur due to differences in specific linear modules. The primary goal of MoM is to enable sparse activation and expand the memory capacity of linear sequence models. To achieve this, we applied sparse activation only to the key and value projections, resulting in a minor increase in activated parameters. This small increase is well justified given the substantial improvements in performance. We’ve also conducted comparisons with models that traditionally use larger parameter counts to expand memory capacity. A detailed discussion on fairness is provided in Appendix F.
4.2 Main Results
4.2.1 Recall-intensive Tasks
Linear sequence models, due to their limited memory capacity, often exhibit a significant performance gap compared to Transformer models, especially in recall-intensive tasks where extensive context is crucial. These tasks highlight notable performance differences among various linear models, making them a more accurate benchmark for evaluating a linear model’s capabilities in handling contextual information.
To thoroughly assess our model’s proficiency in such scenarios, we test six recall-intensive tasks following Arora et al. (2024): FDA Arora et al. (2023), SWDE Arora et al. (2023); Lockard et al. (2019), SQuAD Rajpurkar et al. (2018), NQ Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017) and Drop Dua et al. (2019). These tasks are designed to challenge a model’s ability to perform context-based retrieval and comprehension.
As shown in Table 3.2.2, our proposed approach, benefiting from increased memory capacity and memory mixing mechanism, achieves significant improvements over other linear sequence models. Specifically, our model effectively narrows the performance gap with Transformer models. This improvement underscores the advantage of our method in capturing and utilizing long-range dependencies, thereby enhancing performance on tasks that require extensive contextual understanding.
4.2.2 Commonsense Reasoning Tasks
As shown in Table 3.3, we report the language modeling perplexity and zero-shot performance of commonsense reasoning tasks following Zhang et al. (2024) which includes WikiText Merity et al. (2016), LAMBADA Paperno et al. (2016), ARC-easy, ARC-challenge Clark et al. (2018), HellaSwag Zellers et al. (2019), PiQA Bisk et al. (2020) and WinoGrande Sakaguchi et al. (2019). The evaluation results are based on the lm-evaluation-harness Gao et al. (2024).
Experimental results show that MoM outperforms other linear models and surpassed the Transformer model as well.
Model Sum FS Syn Code Avg.
RetNet 6.30 15.76 2.64 40.52 13.61
HGRN2 6.51 15.50 2.61 40.11 13.02
GSA 7.75 20.29 1.92 42.83 14.61
Gated DeltaNet 7.14 18.00 2.10 41.52 13.98
MoM 6.89 21.26 2.63 47.79 15.64
Table 4: LongBench Benchmark Results. All evaluations were done using OpenCompass. Note: Sum = Summarization, FS = Few-shot, Syn = Synthetic.
Model FDA SWDE SQUAD NQ TriviaQA Drop Avg.
Transformer++ 46.14 25.87 33.22 18.94 45.97 20.03 31.70
MoM 22.98 29.90 29.69 16.60 48.82 20.99 28.16
MoM Hybrid 58.13 44.05 35.71 20.18 48.10 20.60 37.80
Model
ARC-e
$\uparrow$
ARC-c
$\mathrm{acc_{n}}$$\uparrow$
Hella.
$\mathrm{acc_{n}}$$\uparrow$
Lamb.
$\mathrm{acc}$$\uparrow$
PIQA
$\mathrm{acc}$$\uparrow$
Wino.
$\mathrm{acc}$$\uparrow$
Avg.
Transformer++ 44.91 25.94 34.95 26.90 64.31 51.07 41.35
MoM 44.65 24.74 36.54 27.93 66.16 51.78 41.97
MoM Hybrid 46.55 24.49 36.45 28.86 65.51 52.41 42.38
Table 5: Hybrid Model Performance. The hybrid model integrates 1 Transformer layer after every 7 MoM layers, resulting in only 3 Transformer layers across a total of 24 layers.
4.2.3 Long Context Tasks
Assessing performance on long-context tasks is crucial for linear models, as it reflects their ability to handle long-range dependencies effectively. We evaluated our model’s comprehension of long contexts using the LongBench benchmark Bai et al. (2024); Contributors (2023). In Table 4, we present the average results across various categories, including summarization, few-shot learning, synthetic tasks, and code completion, along with the overall mean across all tasks. The complete detailed results are provided in Appendix 10.
Model Params
ARC-e
$\uparrow$
ARC-c
$\mathrm{acc_{n}}$$\uparrow$
Hella.
$\mathrm{acc_{n}}$$\uparrow$
Lamb.
$\mathrm{acc}$$\uparrow$
PIQA
$\mathrm{acc}$$\uparrow$
Wino.
$\mathrm{acc}$$\uparrow$
Avg.
GLA expanded 425M 42.34 22.95 34.56 20.45 63.00 50.12 38.90
GLA MoM 395M 42.85 24.15 36.60 23.23 63.22 49.88 39.99
Gated DeltaNet expanded 550M 43.60 24.66 37.80 26.90 64.47 50.51 41.32
Gated DeltaNet MoM  [^4] M 44.65 24.74 36.54 27.93 66.16 51.78 41.97
Model Params FDA SWDE SQUAD NQ TriviaQA Drop Avg.
GLA expanded 425M 15.08 20.15 28.28 13.30 41.65 18.74 22.87
GLA MoM 395M 9.90 21.65 29.36 14.16 45.20 20.89 23.53
Gated DeltaNet expanded 550M 18.26 24.27 30.03 17.74 48.34 19.26 26.32
Gated DeltaNet MoM  [^4] M 22.98 29.90 29.69 16.60 48.82 20.99 28.16
Table 6: Comparison of Mixture of Memories and Single Memory Expanded. We constructed MoM models using different memory update mechanisms. Separate memory segments yielded better performance compared to simply increasing the memory capacity of a single memory.
4.2.4 The Hybrid of MoM and Transformer
We delve deeper into the hybridization of MoM and Transformer layers by integrating 1 Transformer layer after every 7 MoM layers, resulting in only 3 Transformer layers across a total of 24 layers. The performance on commonsense reasoning and recall-intensive tasks is presented in the table 5. MoM-Hybrid demonstrates significantly improved results compared to Transformer models, despite using only 3 layers of global attention.
4.2.5 Mixed Memory vs. Single Memory
To validate the effectiveness of our mixed memory mechanism, we compare our MoM model with mixed memories to a baseline model that uses an expanded single memory with the same activated memory capacity. We adopt the same memory update method as existing linear models and extend it within our MoM framework. For comparison, we employed the commonly used method of expanding the single memory by expanding the dimension of $v$ to match the total size of all activated memories in the MoM model. Note that while our MoM model activates multiple memories during retrieval, the mixing process effectively consolidates these into a memory capacity equivalent to a single memory. In contrast, the expanded single memory approach increases the entire retrieval capacity which leads to a significant increase in the number of parameters. We evaluate their performance on common-sense reasoning tasks and recall-intensive tasks in Table 6.
The experimental results demonstrated that using multiple mixed memories leads to a greater improvement than simply expanding the capacity of a single memory. This confirms that mixed memory can effectively reduce interference from different inputs. Assigning inputs specifically to different memories, combined with the use of a forget gate, proves to be a more effective approach for reducing interference than relying solely on a forget gate.
Figure 2: Inference Efficiency of MoM. We demonstrate the inference (decoding) time and GPU memory consumption required to generate 1K tokens at specific sequence lengths.
4.2.6 Efficiency
We compare the inference speed and memory usage of MoM and Transformer++ with flash attention in Fig 2. Our analysis demonstrates that MoM exhibits linear complexity, showcasing significant advantages over the Transformer model when handling long sequences. Specifically, MoM’s efficient memory update mechanisms allow it to process longer inputs with reduced computational overhead, positioning it as a more scalable solution for large-scale natural language processing tasks.
4.2.7 Length Extrapolation
We pretrained the models on the Slimpajama dataset with a 2K context length and conducted extrapolation experiments on various lengths using the FinewebPenedo et al. (2024) dataset. We extended the length to 32K to calculate perplexity (ppl). As shown in Fig 3, the Transformer model experienced a significant increase in ppl due to its poor extrapolation capability. Among the linear models, MoM achieved the best results.
Figure 3: Length Extrapolation. We extrapolated models trained on 2K sequences to a length of 32K for perplexity (ppl) evaluation. The ppl of Transformer++ rises sharply from 13.4 at 2k to beyond the display range, reaching 2271.3 at 32k.
Figure 4: Memory Load Balance Analysis. Token Routing Distribution Across Layers and Memories with and without Aux Loss.
4.2.8 Memory Analysis
Memory Load Balance Analysis. To evaluate whether each memory segment in MoM is effectively balanced during inference on downstream tasks, we analyzed the number of tokens routed to each layer using around 300k tokens from the ARC-easy benchmark. We visualized the results with and without auxiliary loss in Fig 4 with heatmaps. Due to the adoption of auxiliary loss, the memory segments in each layer are almost uniformly routed and activated.
Functions of Different Memories. We analyzed the memory routing within the model and identified the categories of tokens processed by each of the four memories. By examining an intermediate layer, we observed that, unlike traditional MoE in FFN layers where experts often lack clear specialization and focus on syntactic punctuationJiang et al. (2023), MoM displays distinct specialization among its memories. This finding suggests potential for exploring an MoE architecture where each memory serves a specific role. These distinctions among different memories are presented in Table 7.
4.2.9 Ablation Study
To evaluate the influence of memory hyperparameters in our MoM model, we conducted ablation studies using a 340 million parameter model trained on 15 billion tokens. Our investigation primarily focused on the impact of the number of memories, the number of activations, and the use of shared memory on model performance. The results of these experiments are presented in Table 8.
Memory Index Type of Tokens Potential Function
1 Basic nouns/verbs/prepositions Simplify semantic information/causal logic
2 Proper nouns/scientific terms Specialized knowledge
3 Technical terms/adjectives Detailed memory
4 Interrogative words/incomplete nouns Fragmented/open information
Table 7: Functions of Different Memories. By analyzing the types of tokens routed to the model’s intermediate layer, we observed a degree of specialization among the different memories.
Acc $\uparrow$
Number of Memories and Activations
(all cases include shared memory by default)
Memories Activated
2 1
25.05
3 2
27.31
4 2
27.59
8 2
25.56
Shared Memory
w/ shared memory 27.59
w/o shared memory 26.06
Table 8: Ablation Study. We performed ablation studies on the number of memories and the use of shared memory. The table presents the average results across all recall-intensive tasks.
5 Conclusion
In this paper, we propose Mixture-of-Memories (MoM), a novel architecture that enhances memory capacity and eliminates memory interference, inspired by the brain’s mechanisms for long-term memory retention. By leveraging multiple independent memory states, MoM significantly improves performance on recall-intensive tasks while maintaining the efficiency advantages of linear-time training and constant-memory inference. Instead of simply discarding tokens as done in gating mechanisms, our memory separation paradigm provides a more effective way to preserve sequence information. Our experimental results demonstrate that MoM outperforms existing linear sequence modeling methods, particularly on tasks requiring strong recall, and achieves performance comparable to Transformer models. This makes MoM a promising approach for applications need strong efficiency and recall-intensive performance, paving the way for sparse solutions in sequence modeling.
6 Acknowledgments
This work is supported by the Shanghai Artificial Intelligence Laboratory.
Appendix A Related Work
Linear Recurrent Models
Linear recurrent models, comprising linear attention, linear RNNs, state-space models (SSMs), have garnered significant research interests Qin et al. (2023b). The advancement of SSMs began with the pioneering work on S4 Gu et al. (2022), which was later optimized through a diagonalized version Gupta et al. (2022). Despite their strong performance on the LRA benchmark, these models have faced challenges in language modeling mainly because they rely solely on data-independent processes. As research progressed, constant forgetting gates were introduced, helping to alleviate some interference issues by uniformly managing memory decay Sun et al. (2023); Gu and Dao (2024). The next breakthrough involved data-dependent forget gates. These allowed models to dynamically adjust memory updates based on the input data, significantly enhancing performance across various tasks Qin et al. (2024c, d); Yang et al. (2023); Zhang et al. (2024); Yang et al. (2024); Qin et al. (2024a). Sequence parallelism techniques are well adapted on linear recurrent models Sun et al. (2024a, 2025) for efficient long context training. There have also been recent advancements in scaling law and test-time regression optimization Shen et al. (2024); Sun et al. (2024c); Behrouz et al. (2024).
Building on these advancements, our MoM model incorporates data-dependent mechanisms that selectively update memory. By efficiently managing interference through tailored memory updates and leveraging increased memory capacity, MoM represents a further evolution, improving model expressiveness and performance.
Mixture-of-Experts
Mixture-of-Experts (MoE) is a technique designed to enhance the capacity of deep neural networks while maintaining computational efficiency Fedus et al. (2022); Rajbhandari et al. (2020); Lepikhin et al. (2020); Tang et al. (2023); Zhu et al. (2024); Qu et al. (2024). MoE achieves this by activating a subset of parameters, known as "experts", for each input, which reduces the computational costs. Shazeer first integrated MoE into LSTM layers Shazeer et al. (2017). The Switch Transformer Fedus et al. (2022) refined this approach by simplifying the gating mechanism to select only one expert per input. Gshard Lepikhin et al. (2020) further advanced this by using a top-2 expert routing strategy to improve performance. Recent MoE models, such as Deepseek-MoE Dai et al. (2024), introduce shared experts to capture and consolidate common knowledge across different contexts, while designing fine-grained experts to increase combinatorial flexibility.
Appendix B Comparison between MoM and MoE
While our approach to implementing the Mixture-of-Memories (MoM) draws inspiration from the Mixture-of-Experts (MoE) framework, there are significant differences that distinguish our method from traditional MoE implementations.
  • Purpose: The MoE was introduced to scale up the number of parameters without significantly increasing computational resources. It address the limitations of dense models in scaling both parameters and computational demands through sparse activation. However, MoM is designed to expand the memory capacity of linear attention models while preserving their linear time complexity. By sparsely activating memories and using weighed summation to create a mixed memory, MoM effectively address the challenge of forgetting historical information in linear attention. Moreover, by separating the memory into distinct states, MoM reduces interference between different pieces of information.
  • Structure: In conventional MoE, each expert is a separate neural network within the feedforward network (FFN) layer Team (2024). In contrast, in MoM, each memory is an RNN state with unique key-value projection weights to generate different key-value pairs. MoE operates during the channel mixing phase, where each token is processed independently by selected experts. On the other hand, MoM functions during the token mixing phase, where each memory processes different segments of the sequence, preserving inter-token relationships.
Appendix C Dataset and Benchmark
We pretrained model on SlimPajama dataset. For 340M models, we train on a sample of 15B tokens. For 1.3B models, we train on a sample of 100B tokens.
  • SlimPajama Soboleva et al. (2023) is a high-quality, optimized subset of the RedPajama dataset, designed for large-scale language model training. It includes diverse text sources such as Common Crawl, Wikipedia, books, and GitHub code, with a primary focus on English. The dataset is cleaned, deduplicated, and optimized for efficiency and performance.
For the benchmark, we tested on these tasks:
  • WikiText Merity et al. (2016): A dataset consisting of high-quality Wikipedia articles, primarily in English, designed for language modeling tasks with 62 test samples. The text covers a wide range of topics, including history, science, and culture, and is authored by Wikipedia contributors, who come from diverse demographic backgrounds.
  • LAMBADA Paperno et al. (2016): An English-language dataset for evaluating contextual understanding in language models with 5153 test samples. It consists of narrative texts sourced from books, requiring models to predict the final word of a passage. The data reflects a mix of literary styles and author demographics.
  • ARC-Easy & ARC-Challenge Clark et al. (2018): A set of multiple-choice science questions in English, sourced from standardized exams and educational materials with 2376 and 1172 test samples. The dataset represents the domain of elementary and high school science, with questions authored by educators and test designers. ARC-Easy includes straightforward questions, while ARC-Challenge contains more difficult ones that require advanced reasoning.
  • HellaSwag Zellers et al. (2019): An English-language dataset designed for commonsense reasoning, where models must choose the most plausible continuation of a sentence. The text is derived from activity descriptions (e.g., WikiHow), covering everyday scenarios. The dataset was constructed adversarially to be challenging for language models. It has 10003 test samples.
  • PiQA Bisk et al. (2020): A dataset focused on physical commonsense reasoning in English with 3084 test samples. The text consists of everyday tasks and scenarios, requiring models to determine the most practical way to perform an action. The data is sourced from crowdsourced descriptions, reflecting a broad range of common human experiences.
  • WinoGrande Sakaguchi et al. (2019): A large-scale English dataset for commonsense reasoning, based on the Winograd Schema Challenge with 1267 test samples. It tests pronoun resolution in ambiguous contexts, with sentences sourced and refined through crowdsourcing. The dataset aims to reduce annotation biases by diversifying sentence structures and topics.
  • FDA Arora et al. (2024, 2023): An English-language dataset includes 100 randomly sampled PDF files of FDA 510(k) pre-market notification submissions, each with 16 manually annotated attributes like device classification and predicate device codes. It has 1102 test samples.
  • SWDE Arora et al. (2024, 2023); Lockard et al. (2019): An English dataset dataset is used for evaluating information extraction systems. It comprises data from 8 movie websites (e.g., IMDB, Rotten Tomatoes) and 5 university websites (e.g., US News), with each site containing 1063 to 2000 annotated web pages and $1$ test samples.
  • SQuAD Rajpurkar et al. (2018): A reading comprehension dataset with English questions based on Wikipedia articles with 2984 test samples. Answers are text segments from the articles or the questions can be unanswerable. SQuAD2.0 adds over 50,000 unanswerable questions to the original 100,000 from SQuAD1.1, requiring systems to both answer questions and identify when no answer is supported.
  • NQ Kwiatkowski et al. (2019): A large-scale English dataset for open-domain question answering, where questions are sourced from real Google search queries. Answers are extracted from Wikipedia articles, reflecting general knowledge across various domains. It has 3157 test samples
  • TriviaQA Joshi et al. (2017): An English-language question-answering dataset consisting of trivia-style questions with 1688 test samples. The dataset includes both human-generated questions and web-sourced evidence, covering a broad range of topics such as history, science, and entertainment.
  • DROP Dua et al. (2019): A reading comprehension benchmark in English that requires discrete reasoning, such as arithmetic and logical operations, over passages. The dataset is sourced from Wikipedia and challenges models to go beyond simple span extraction by performing multi-step reasoning. It has 2087 test samples.
  • LongBench Bai et al. (2024) is a bilingual (Chinese and English) benchmark for evaluating long-context understanding in large language models. It covers 21 tasks across six categories, including QA, summarization, and code completion with 4750 test samples.
All datasets used in this work are publicly available and have been released by their original creators, who are responsible for ensuring privacy protection. These datasets are used in accordance with their respective licenses and intended purposes. No modifications or derivative datasets have been created.
Appendix D Experiments Details
Our experiments were conducted using 32 NVIDIA A800 GPUs. Training the 340M parameter model required approximately 10 hours, while the 1.3B parameter model took around 6 days. All models were trained and evaluated using a fixed random seed of 42 to ensure reproducibility. The reported results are based on a single run, without aggregation over multiple random seeds.
Appendix E Training Loss Comparison
To further assess the learning efficiency of MoM, we compared the training loss curves of MoM with those of other baseline models. As depicted in Figure 5, MoM consistently maintains the lowest loss throughout the entire training phase. Even as training nears convergence, MoM continues to exhibit a clear advantage over other methods.
Figure 5: Training Loss. Loss curves for training 340M models on 15B tokens with a fixed random seed of 42.
Appendix F Fairness
To enhance memory capacity, MoM applies sparse activation to the key and value projections. Although these projections constitute a small portion of the overall model parameters, this inevitably increases the parameter count. Due to differing linear model structures, aligning both parameter count and memory capacity exactly is challenging. Thus, to ensure fairness, we conduct comparisons from two perspectives: equal activated parameter count and equal memory capacity.
F.1 Equal Activated Parameter Count
To ensure a fair comparison of parameters, we reduced the MLP hidden ratio to 2 and retrained the MoM model using the same training configurations as in Section 4.1. Both MoM and Gated Deltanet were set with 400M activated parameters. Although the smaller hidden ratio might impact the model’s commonsense knowledge, we tested on commonsense reasoning tasks and recall-intensive benchmarks. MoM consistently outperformed Gated Deltanet in both tests, further validating the effectiveness of the MoM approach. The results are presented in Table 9
Model Params
ARC-e
$\uparrow$
ARC-c
$\mathrm{acc_{n}}$$\uparrow$
Hella.
$\mathrm{acc_{n}}$$\uparrow$
Lamb.
$\mathrm{acc}$$\uparrow$
PIQA
$\mathrm{acc}$$\uparrow$
Wino.
$\mathrm{acc}$$\uparrow$
Avg.
Gated DeltaNet 400M 46.04 23.55 35.18 27.01 66.05 50.83 41.44
MoM 400M 47.10 23.72 35.43 26.88 64.64 51.22 41.50
Model Params FDA SWDE SQUAD NQ TriviaQA Drop Avg.
Gated DeltaNet 400M 20.53 23.24 28.55 14.98 44.91 16.48 24.78
MoM 400M 24.16 25.59 29.46 15.36 46.15 18.35 26.51
Table 9: Comparison with the Same Activated Parameters. MoM and Gated DeltaNet with 400M activated parameters are tested.
F.2 Equal Memory Capacity
To ensure a fair comparison of memory capacity, we also compared the single extended memory model with the MoM model. Notably, the single extended memory model has more parameters than the activated parameters in the MoM due to the extension of the $v$ dimension. MoM expands memory more elegantly and significantly outperforms in both recall and commonsense tasks. This comparison result is presented in Table 6.
Model SQA MQA Sum FS Syn Code Zh-Avg En-Avg Avg.
RetNet 9.23 7.23 6.3 15.76 2.64 40.52 15.44 13.5 13.61
HGRN2 7.38 6.02 6.51 15.5 2.61 40.11 14.28 13.12 13.02
GSA 8.21 6.63 7.75 20.29 1.92 42.83 15.06 15.2 14.61
Gated DeltaNet 8.52 6.61 7.14 18 2.1 41.52 14.19 14.63 13.98
MoM 8.14 7.11 6.89 21.26 2.63 47.79 17.33 15.71 15.64
Table 10: Complete Results of LongBench. SQA: Single-doc QA, MQA: Multi-doc QA, Sum: Summarization, FS: Few-shot learning, Syn: Synthetic
FDA SWDE SQUAD NQ TriviaQA Drop Avg.
Number of Memories
Memories Activated
2 1
20.53 23.15 29.56 13.72 43.96 19.36 25.05
3 2
23.71 29.52 28.69 15.55 46.09 20.32 27.31
4 2
30.79 26.05 29.63 13.84 44.79 20.41 27.59
8 2
19.71 24.27 30.13 15.68 46.15 17.39 25.56
Shared Memory
w/ shared memory 30.79 26.05 29.63 13.84 44.79 20.41 27.59
w/o shared memory 25.61 27.09 28.35 14.32 41.94 19.02 26.06
Table 11: Complete Ablation Study Results. The complete results on recall-intensive tasks.
Wiki.
$\downarrow$
Lamb.
$\downarrow$
ARC-e
$\uparrow$
ARC-c
$\mathrm{acc_{n}}$$\uparrow$
Hella.
$\mathrm{acc_{n}}$$\uparrow$
Lamb.
$\mathrm{acc}$$\uparrow$
PIQA
$\mathrm{acc}$$\uparrow$
Wino.
$\mathrm{acc}$$\uparrow$
Avg.
Number of Memories
Memories Activated
2 1
26.84 66.26 46.25 23.21 35.54 25.40 64.69 51.30 41.07
3 2
25.86 52.78 45.62 23.46 36.25 28.84 65.13 53.67 42.16
4 2
26.00 51.25 46.13 24.15 35.91 28.26 65.61 52.57 42.10
8 2
26.40 52.30 45.66 24.74 35.93 28.39 64.91 49.49 41.52
Shared Memory
w/ shared memory 26.00 51.25 46.13 24.15 35.91 28.26 65.61 52.57 42.10
w/o shared memory 27.74 85.22 45.33 24.15 34.91 23.52 64.09 50.28 40.38
Table 12: Complete Ablation Study Results. The complete results on common-sense reasoning tasks.