### v4

# MoM: Linear Sequence Modeling with Mixture-of-Memories

Abstract
Linear sequence modeling methods, such as linear attention, state space modeling, and linear RNNs, offer significant efficiency improvements by reducing the complexity of training and inference. However, these methods typically compress the entire input sequence into a single fixed-size memory state, which leads to suboptimal performance on recall-intensive tasks. To address this limitation, we introduce a novel architecture called Mixture-of-Memories (MoM). MoM utilizes multiple independent memory states, with a router network directing input tokens to specific memory states. This approach greatly enhances the overall memory capacity while minimizing memory interference. MoM serves as a general framework that can be seamlessly combined with diverse memory update mechanisms across linear models. As a result, MoM performs exceptionally well on recall-intensive tasks, surpassing existing linear sequence modeling techniques. Despite incorporating multiple memory states, the computation of each memory state remains linear in complexity, allowing MoM to retain the linear-complexity advantage during training, while constant-complexity during inference. Our experimental results show that MoM outperforms current linear sequence models on downstream language tasks, particularly recall-intensive tasks, and even achieves performance comparable to Transformer models. The code is released at https://github.com/OpenSparseLLMs/MoM and is also released as a part of https://github.com/OpenSparseLLMs/Linear-MoE.
* Interns at Shanghai AI Laboratory; 🖂 Corresponding Authors: Weigao Sun (sunweigao@outlook.com) and Yu Cheng (chengyu@cse.cuhk.edu.hk); † Project Lead.
”The enemy of memory isn’t time; it’s other memories.”
– David Eagleman
1 Introduction
Attention mechanisms have made significant contributions to the field of artificial intelligence, advancing various modalities such as language, vision, audio, video, graphs, and even time series (achiam2023gpt; team2023internlm). The Transformer (vaswani2017attention), known for its ability to capture long-range dependencies, has become a foundational architecture in this space. However, traditional Transformers encounter computational challenges due to their quadratic time complexity, $O(n^{2})$, with respect to sequence length $n$, making it difficult to scale to long sequences. To overcome this limitation, several linear sequence modeling methods have been proposed, including linear attention (katharopoulos2020transformers; qin2023transnormerllm; li2025minimax), state space modeling (gu2024mambalineartimesequencemodeling; dao2024transformers), and linear RNNs (peng2024eagle; qin2024hgrn2), which offer $O(n)$ training complexity and $O(1)$ inference complexity. These approaches often reduce the input sequence to a fixed-size hidden space, collapsing the information into a single “memory state”. While these methods enhance efficiency, they face two main challenges: limited memory capacity and memory interference. When new information overwrites the single fixed-size memory state, previously stored representations may degrade, which negatively impacts its long-term memory performance on recall-intensive tasks.
We argue that the strong performance of Transformer models on recall-intensive tasks arises from their ability to avoid memory interference by maintaining independent key-value caches for each token, thus offering virtually unlimited memory capacity. In contrast, linear sequence modeling relies on extreme compression, consolidating all the input information into a single fixed-size memory state (katharopoulos2020transformers; dao2024transformers). This approach results in limited memory capacity and inherently leads to memory interference issues.
[…]
Inspired by these biological insights, we propose a new architecture called Mixture-of-Memories (MoM), which aims to strike a balance between the explicit token representations in Transformers and the extreme compression found in earlier linear sequence modeling methods. MoM employs multiple independent memory states, with a router network that directs input tokens to specific memory states. The input sequence is divided into a predefined number of subsequences (phase-specific neural assemblies), which are processed in parallel and fed into the corresponding memory projections (dentate microcircuits) to generate key-value pairs. As the linear sequence modeling layer processes each subsequence using an RNN-like update mechanism, it produces multiple memory states that capture different aspects of the input sequence. The final output is computed as a weighted sum of these memories, which we refer to as the mixture-of-memories. This approach expands memory capacity and eliminates memory interference, enabling MoM to significantly outperform existing linear sequence models that rely on a single fixed-size memory state.
Our contributions can be summarized as follows:
  • We present MoM, an architecture that incorporates multiple independent memory states, significantly enhancing memory capacity and eliminating memory interference, while retaining the efficiency benefits of linear-time training and constant-memory inference.
  • Distinct with existing gating mechanisms, MoM is a new paradigm to reduce memory interference by separating the memory states. The overall design is broadly compatible with diverse linear sequence modeling methods, making it a straightforward and effective approach to boost task performance.
  • Through empirical evaluation, we show that MoM outperforms strong linear sequence modeling baselines across a variety of language tasks, particularly on recall-intensive tasks. MoM even achieves performance on par with Transformer models, a feat that current linear sequence modeling methods struggle to match.
2 Preliminary
For notations in this work, we use bold lower-case letters for row vectors (e.g., $\bm{q}_{t},\bm{k}_{t}$), bold upper-case letters for matrices (e.g., $\bm{Q},\bm{K}$) and the identical letters represent a row in the matrix, e.g., $\bm{q}_{t}$ is the $t$-th row of $\bm{Q}$.
Linear Attention
To reduce the time complexity of Transformer attention, various optimization techniques have been proposed. Linear Transformers (katharopoulos2020transformers) replace the softmax attention mechanism with dot-product of feature maps $\phi(\cdot)$:
$\bm{o}_{t}=\frac{\sum_{i=1}^{n}\phi(\bm{q}_{t})\phi(\bm{k}_{i})^{T}\bm{v}_{i}}{\sum_{i=1}^{n}\phi(\bm{q}_{t})\phi(\bm{k}_{i})^{T}},$ (1)
where $\bm{q}_{t},\bm{k}_{t},\bm{v}_{t}\in\mathbb{R}^{d}$. While the presence of the denominator may lead to numerical instability (qin2024lightning) and the feature map can utilize an identity function, which we omit for simplicity. In perspective of memory, the formulation can also be written in a recurrent format:
$\bm{M}_{t}=\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t},\quad\bm{o}_{t}=\bm{q}_{t}\bm{M}_{t}.$ (2)
This indicates that linear attention can function as a linear recurrent layer with a matrix-valued hidden state $\bm{M}$ which we refer to as memory sate and the output is generated by querying the memory state $\bm{M}$. This represents the ultimate compression of sequence information, condensing the entire sequence into a single memory state.
Building on the foundational concepts of linear attention and memory perspective, some recent advancements have focused on optimizing memory structure, including gated updates (yang2023gated; qin2024hierarchically; qin2024hgrn2) and memory capacity expansion (peng2024eagle; qin2024hgrn2).
3 Method
3.1 Motivation
Linear sequence models compress the entire sequence data into a fixed-size memory state. Despite numerous efforts to minimize information loss, such as introducing gating mechanisms and employing more precise control over memory modifications (orvieto2023resurrecting; de2024griffin; beck2024xlstm; yang2023gated; zhang2024gated), some degradation in this compression process is inevitable. Expanding the memory capacity has been shown to mitigate this issue to some extent, with studies indicating that increasing memory capacity can enhance model performance (qin2024hgrn2; peng2024eagle).
However, previous approaches that simply increased the size of the RNN state, essentially expanding a single memory state, struggled to capture the full spectrum of information within an entire sequence. We propose that this difficulty arises because sequence information is often multifaceted, and a single, expanded memory may not be capable of simultaneously capturing multiple aspects of the data. Inputs that introduce new or orthogonal information may interfere with existing memory content when using a shared memory. Rather than discarding these inputs through gating mechanisms or overwriting the existing memory state, it may be more effective to consider alternative strategies that allow for the preservation of diverse information without interference.
3.2 MoM: Mixture-of-Memories
Figure 1: MoM Architecture. Each input token selectively activates and updates $K$ memory states, leaving non-activated memory states unchanged to avoid interference from current input. Additionally, we introduce a continuously activated shared memory. This figure presents the basic memory update mechanism; other mechanisms involving gating or more complex updates follow a similar approach.
To address the challenge outlined above, we propose a novel approach for encoding multi-item memory such as theta-gamma oscillations (lisman2013theta), and concepts from Mixture-of-Experts (MoE) (shazeer2017outrageously), where different experts handle specific tokens. In this approach, we leverage multiple memory states, each of which is selectively updated by different inputs. This increases the memory capacity and enables the model to retain diverse pieces of information by storing various types of inputs in separate memory states.
In our framework, the memory states function similarly to the experts in MoE. However, instead of relying on completely separate networks, these modules are individual RNN states embedded within a linear recurrent mechanism. This design allows for the isolation of memory updates while concurrently managing distinct types of information. It is important to note that MoM fundamentally differs from traditional MoE, as we will discuss in Appendix B. Figure 1 provides an overview of the MoM architecture. Below, we introduce the structure of the MoM layer and explain how this multi-memory architecture is implemented in the context of linear sequence modeling.
3.2.1 Router
We use a router to assign inputs to different memory states. Utilizing the top-$k$ concept, each token is routed to the top-$k$ memories based on its importance scores. Specifically, we use a simple linear layer to generate these scores for each input token. After applying a softmax function, we select the top-$k$ scores and normalize them.
$\mathbf{scores}_{t}=\mathrm{TopK}(\mathrm{softmax}(\bm{x}_{t}\bm{W}_{g}))\in\mathbb{R}^{k},$ (3)
$\bm{g}_{t}=\frac{\mathbf{scores}_{t}}{\sum\mathbf{scores}_{t}}\in\mathbb{R}^{k},$ (4)
where $\bm{x}_{t}\in\mathbb{R}^{d}$, $k$ is the top-$k$ number, $\bm{W}_{g}\in\mathbb{R}^{d\times M}$ is learnable weight, $\bm{g}_{t}$ is the normalized importance scores of the input $\bm{x}_{t}$.
3.2.2 Linear Recurrent Memory Module
After the router network, the input $\bm{x}_{t}$ is directed to top-$k$ linear recurrent modules, meaning that the top-$k$ memories are activated while the others remain inactive.
Each Memory. For each activated memory, indexed by $m$, we perform the following operation:
Table 1: Memory Update Rules. We demonstrate that several linear sequence models can be viewed as recurrent models in terms of memory updates, where $a_{t},b_{t}\in(0,1)$ are data-dependent scaler, $\bm{a}_{t}$ is data-dependent vector, and $\gamma$ is a data-independent constant.
Method Memory Update Rule Linear Attn $\bm{M}_{t}=\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$ RetNet $\bm{M}_{t}=\gamma\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$ GLA $\bm{M}_{t}=(\bm{a}_{t}^{T}\bm{1})\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$ DeltaNet $\bm{M}_{t}=(\bm{I}-\bm{k}_{t}^{T}\bm{k}_{t})\bm{M}_{t-1}+b_{t}\bm{k}_{t}^{T}\bm{v}_{t}$ G-DeltaNet $\bm{M}_{t}=a_{t}(\bm{I}-\bm{k}_{t}^{T}\bm{k}_{t})\bm{M}_{t-1}+b_{t}\bm{k}_{t}^{T}\bm{v}_{t}$ TTT $\bm{M}_{t}=\bm{M}_{t-1}+b_{t}\nabla l(\bm{M}_{t-1};\bm{k}_{t},\bm{v}_{t})$ Titans $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+b_{t}\nabla_{M}l(\bm{M}_{t-1};\bm{k}_{t},\bm{v}_{t})$ Mamba2 $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+b_{t}\bm{k}_{t}^{T}\bm{v}_{t}$ HGRN2 $\bm{M}_{t}=(\bm{a}_{t}^{T}\bm{1})\bm{M}_{t-1}+(1-\bm{a}_{t})^{T}\bm{v}_{t}$ RWKV6 $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$ RWKV7 $\bm{M}_{t}=(\bm{a}_{t}^{T}\bm{1})\bm{M}_{t-1}+b_{t}\nabla l(\bm{M}_{t-1};\bm{k}_{t},\bm{v}_{t})$
  1. 1.
    Key and Value Projections: We project the input $\bm{x}_{t}$ to $\bm{k}^{m}_{t}$ and $\bm{v}^{m}_{t}$ using $\bm{W}^{m}_{k}$ and $\bm{W}^{m}_{v}$:
    $\bm{k}^{m}_{t}=\bm{x}_{t}\bm{W}^{m}_{k},\bm{v}^{m}_{t}=\bm{x}_{t}\bm{W}^{m}_{v}\in\mathbb{R}^{d},$ (5)
    where $\bm{W}^{m}_{k}$, $\bm{W}^{m}_{v}$ are learnable projection weights for $kv$ of the $m$-th memory module.
  2. 2.
    Memory Update: We update the activated memory state using $\bm{k}^{m}_{t}$, $\bm{v}^{m}_{t}$:
    $\bm{M}^{m}_{t}=\bm{M}^{m}_{t-1}+(\bm{k}^{m}_{t})^{T}\bm{v}^{m}_{t}\in\mathbb{R}^{d\times d}.$ (6)
    The equation above represents the simplest form of memory update for clarity. Our approach is flexible and does not rely on a specific memory update mechanism. To enhance performance, we can incorporate mechanisms such as forget gates (sun2023retentive).
    More generally, our method can be adapted to incorporate various memory update methods proposed in previous work. Detailed descriptions of these methods are provided in Table 1.
Memory Mixing. After updating the activated memory states, we perform a weighted sum of these memory states using the importance scores obtained from Equation(4).
$\tilde{\bm{M}}_{t}=\sum_{m}g_{t}^{(m)}\bm{M}^{m}_{t}\in\mathbb{R}^{d\times d},$ (7)
where $\bm{M}^{m}_{t}$ is one activated memory and $g_{t}^{(m)}$ is the importance score of $\bm{M}^{m}_{t}$.
We then obtain the output of the MoM by applying query vector $\bm{q}_{t}$ to the mixed memory $\tilde{\bm{M}}_{t}$:
$\bm{o}_{t}=\bm{q}_{t}\tilde{\bm{M}}_{t}\in\mathbb{R}^{d}.$ (8)
Finally, the output of the MoM layer is computed by applying an activation function, normalization, and a linear transformation.
Throughout the recurrent process, only a subset of memory states is activated and updated at each time step, while memory states that are not routed remain inactive and unchanged. When the input passes through the key-value projection layer, it generates multiple sets of keys and values that are fed into different memory modules. This design enables the model to maintain multiple memory states, each preserving distinct pieces of information. By aggregating the activated memories into a comprehensive mixed memory by weighted summation, the query can effectively retrieve information from this mixed memory, and generate attention output followed by other layers.
Shared Memory. To enhance our model’s ability to capture long-term dependencies, we introduce a shared memory mechanism. This shared memory has access to the entire sequence information, allowing it to effectively store and retrieve long-term information. By integrating shared memory into our model, we ensure that it can leverage the complete historical context, resulting in significant improvements in performance and robustness.
Figure 2: Hardware-efficient Implementation of MoM. Tokens sharing the same color are routed to the same memory. ① Tokens are first split into groups according to memory routing results, ② then concatenated into a varlen input sequence, ③ processed by the Triton kernel, ④ the outputs are returned, ⑤ split back into their respective memories, and ⑥ finally restored to the original sequence order. For clarity, the illustration shows the top-$1$ routing case, and the qkvqkv projection is omitted.
3.3 Hardware-Efficient Implementation
In the implementation of MoM, mixing memories before query multiplication is equivalent to multiplying each memory by the query and then mixing the results, allowing us to reuse efficient Triton-based operators from prior linear sequence models. We first reorder the sequence tokens according to the routing results so that they follow the memory layout. The reordered tokens are then concatenated with varlen for operator computation, after which the results are aggregated via weighted summation. In this way, MoM’s computation can be effectively reduced to varlen operations, enabling efficient execution. We elaborate on this process below.
Given input tokens $\bm{x}_{b,t}\in\mathbb{R}^{d}$ for batch $b\in\{1,\dots,B\}$ and time step $t\in\{1,\dots,T\}$, each token is routed to one or more memories $m\in\{1,\dots,M\}$ with routing weights $\alpha_{b,t,m}\geq 0$ satisfying $\sum_{m=1}^{M}\alpha_{b,t,m}=1$.
For each $(b,m)$, define the ordered index set
$\mathcal{I}_{b,m}=\bigl(t_{b,m}(1),\dots,t_{b,m}(L_{b,m})\bigr),$
where $t_{b,m}(j)$ is the original sequence index of the $j$-th token assigned to memory $m$, and $L_{b,m}=|\mathcal{I}_{b,m}|$. We index buckets lexicographically by p=(b1)M+mp=(b-1)M+m and define cumulative boundaries
$s_{0}=0,\qquad s_{p}=\sum_{q=1}^{p}L_{q}\quad(p=1,\dots,BM).$
The flattened sequence $\bm{\tilde{X}}$ is obtained by
$\bm{\tilde{x}}_{\,s_{p-1}+j}=\bm{x}_{\,b,\,t_{b,m}(j)},\qquad j=1,\dots,L_{b,m},$
with varlen representation $(\bm{\tilde{X}},\bm{s})$, where $\bm{s}=(s_{0},\dots,s_{BM})$.
For each bucket p=(b1)M+mp=(b-1)M+m, queries share a projection matrix $\bm{W}_{Q}$, while keys and values use memory-specific projections $\bm{W}_{K}^{(m)},\bm{W}_{V}^{(m)}$:
$\bm{\tilde{q}}_{u}=\bm{W}_{Q}\bm{\tilde{x}}_{u},\quad\bm{\tilde{k}}_{u}=\bm{W}_{K}^{(m)}\bm{\tilde{x}}_{u},\quad\bm{\tilde{v}}_{u}=\bm{W}_{V}^{(m)}\bm{\tilde{x}}_{u},\quad u\in\{s_{p-1}+1,\dots,s_{p}\}.$
A memory-specific kernel $\mathcal{F}_{m}$ with parameters $\bm{\theta}^{(m)}$ is applied independently to each segment:
$\bm{o}_{\,s_{p-1}+1:s_{p}}=\mathcal{F}_{m}\!\bigl(\bm{\tilde{q}}_{s_{p-1}+1:s_{p}},\,\bm{\tilde{k}}_{s_{p-1}+1:s_{p}},\,\bm{\tilde{v}}_{s_{p-1}+1:s_{p}};\,\bm{\theta}^{(m)}\bigr).$
Mapping outputs back to the original sequence, the $j$-th token in $\mathcal{I}_{b,m}$ has per-memory output
$\bm{\hat{o}}_{b,t_{b,m}(j),m}=\bm{o}_{\,s_{p-1}+j}.$
Finally, token-level representations are reconstructed by weighted summation:
$\bm{y}_{b,t}=\sum_{m=1}^{M}\alpha_{b,t,m}\,\bm{\hat{o}}_{b,t,m}.$
4 Experiments
4.1 Experimental Setups
Models.  In our experiments, we employ the Gated DeltaNet (yang2024gated) as the memory update mechanism in MoM. The model is configured with four memory states, two of which are activated at each time step, along with a shared memory.
Baselines.  We evaluate MoM against several linear recurrent models and Transformers, including RetNet (sun2023retentive), GLA (yang2023gated), Gated DeltaNet (yang2024gated), and Transformer++ (touvron2023llama), which incorporates Rotary Position Embeddings (su2024roformer) and GLU (shazeer2020glu) into the Transformer architecture. To ensure a fair comparison, we train all baseline models from scratch using the exact same number of tokens.
Training.  We follow the training procedure described by yang2023gated, utilizing the SlimPajama dataset (cerebras2023slimpajama) sampled with 100B tokens and tokenized using the Mistral tokenizer (jiang2023mistral). We train models from scratch with parameter sizes of 380M and 1.3B, respectively. For the 380M models, we train on 15B tokens with a batch size of 0.5M tokens. More detailed training configuration is provided in Appendix C. We utilized publicly available pretrained weights from zhang2024gated with exactly same configuration  [^1] Models marked with an asterisk use open-source pretrained weights with identical training configurations..
Parameter Explanation. We report model sizes using the common shorthand, where “380M” denotes a configuration with 24 layers and hidden size 1024, and “1.3B” denotes 24 layers with hidden size 2048. The main goal of MoM is to expand the memory capacity of linear sequence models through sparse activation. To this end, we apply sparse activation only to the key and value projections, which results in a small increase in activated parameters that is well justified by the performance gains. A detailed discussion on fairness is provided in Appendix G.
4.2 Main Results
Table 2: Results on Recall-Intensive Tasks. All inputs are truncated to a maximum length of 2K tokens. MoM significantly outperforms all other linear models across both model sizes. In the 1.3B model, MoM even achieves performance very close to that of Transformer models.
Scale Model FDA SWDE SQUAD NQ TriviaQA Drop Avg.
Avg.
(no FDA)
380M Params 15B Tokens L=24, d=1024 Transformer++ 46.14 25.87 33.22 18.94 45.97 20.03 31.70 28.81
RetNet 5.90 9.28 22.41 6.91 40.05 18.59 17.19 19.45
HGRN2 11.53 17.34 24.08 12.67 43.84 17.35 21.14 23.06
GLA 11.26 16.78 27.85 12.77 43.90 17.68 21.71 23.80
GSA 6.36 16.87 21.90 14.60 42.18 16.72 19.77 22.45
Gated DeltaNet 20.53 23.24 28.55 14.98 44.91 16.48 24.78 25.63
MoM 22.98 29.90 29.69 16.60 48.82 20.99 28.16 29.20
1.3B Params 100B Tokens L=24, d=2048 Transformer++ 44.32 32.43 42.59 24.49 58.47 21.56 37.31 35.91
RetNet 13.62 22.59 33.46 15.43 53.79 19.79 26.45 29.01
HGRN2 12.35 23.24 33.19 19.10 55.27 19.65 27.13 30.09
GLA 27.61 30.93 35.04 22.27 56.28 19.45 31.93 32.79
GSA 23.25 32.80 35.57 22.96 57.05 20.65 32.05 33.81
Gated DeltaNet 30.25 27.65 34.06 23.22 58.23 20.36 32.30 32.70
MoM 41.14 34.30 37.08 24.11 58.59 21.03 36.04 35.02
4.2.1 Recall-intensive Tasks
Linear sequence models, due to their limited memory capacity, often exhibit a significant performance gap compared to Transformer models, especially in recall-intensive tasks where extensive context is crucial. These tasks highlight notable performance differences among various linear models, making them a more accurate benchmark for evaluating a linear model’s capabilities in handling contextual information.
To thoroughly assess our model’s proficiency in such scenarios, we test six recall-intensive tasks following arora2024simple: FDA (arora2023language), SWDE (arora2023language; lockard-etal-2019-openceres), SQuAD (rajpurkar2018know), NQ (kwiatkowski2019natural), TriviaQA (joshi2017triviaqa) and Drop (dua2019drop). These tasks are designed to challenge a model’s ability to perform context-based retrieval and comprehension.
As shown in Table 2, our proposed approach, benefiting from increased memory capacity and memory mixing mechanism, achieves significant improvements over other linear sequence models. Specifically, our model effectively narrows the performance gap with Transformer models. This improvement underscores the advantage of our method in capturing and utilizing long-range dependencies, thereby enhancing performance on tasks that require extensive contextual understanding.
Table 3: LongBench Benchmark Results. Note: Sum = Summarization, FS = Few-shot, Syn = Synthetic.
Model Sum FS Syn Code Avg. RetNet 6.30 15.76 2.64 40.52 13.61 HGRN2 6.51 15.50 2.61 40.11 13.02 GSA 7.75 20.29 1.92 42.83 14.61 Gated DeltaNet 7.14 18.00 2.10 41.52 13.98 MoM 6.89 21.26 2.63 47.79 15.64
4.2.2 Long Context Tasks
Assessing performance on long-context tasks is crucial for linear models, as it reflects their ability to handle long-range dependencies effectively. We evaluated our model’s comprehension of long contexts using the LongBench benchmark (bai2024longbenchbilingualmultitaskbenchmark; 2023opencompass). In Table 3, we present the average results across various categories, including summarization, few-shot learning, synthetic tasks, and code completion, along with the overall mean across all tasks. The complete detailed results are provided in Appendix I.
Table 4: Comparison of Mixture of Memories and Single Memory Expanded. We constructed MoM models using different memory update mechanisms. Separate memory segments yielded better performance compared to simply increasing the memory capacity of a single memory.
Model Params
ARC-e
[…]
ARC-c
$\mathrm{acc_{n}}\uparrow$
Hella.
$\mathrm{acc_{n}}\uparrow$
Lamb.
$\mathrm{acc}\uparrow$
PIQA
$\mathrm{acc}\uparrow$
Wino.
$\mathrm{acc}\uparrow$
Avg.
GLA expanded 425M 42.34 22.95 34.56 20.45 63.00 50.12 38.90
GLA MoM 395M 42.85 24.15 36.60 23.23 63.22 49.88 39.99
Gated DeltaNet expanded 550M 43.60 24.66 37.80 26.90 64.47 50.51 41.32
Gated DeltaNet MoM  [^4] M 44.65 24.74 36.54 27.93 66.16 51.78 41.97
Model Params FDA SWDE SQUAD NQ TriviaQA Drop Avg.
GLA expanded 425M 15.08 20.15 28.28 13.30 41.65 18.74 22.87
GLA MoM 395M 9.90 21.65 29.36 14.16 45.20 20.89 23.53
Gated DeltaNet expanded 550M 18.26 24.27 30.03 17.74 48.34 19.26 26.32
Gated DeltaNet MoM  [^4] M 22.98 29.90 29.69 16.60 48.82 20.99 28.16
4.2.3 Mixed Memory vs. Single Memory
To validate the effectiveness of our mixed memory mechanism, we compare our MoM model with mixed memories to a baseline model that uses an expanded single memory with the same activated memory capacity. We adopt the same memory update method as existing linear models and extend it within our MoM framework. For comparison, we employed the commonly used method of expanding the single memory by expanding the dimension of $\bm{v}$ to match the total size of all activated memories in the MoM model. We evaluate their performance on common-sense reasoning tasks and recall-intensive tasks in Table 4.
The experimental results demonstrated that using multiple mixed memories leads to a greater improvement than simply expanding the capacity of a single memory with less parameters. This confirms that mixed memory can effectively reduce interference from different inputs. Assigning inputs specifically to different memories, combined with the use of a forget gate, proves to be a more effective approach for reducing interference than relying solely on a forget gate.
4.2.4 Efficiency
We compare the inference speed and memory usage of MoM and Transformer++ with flash attention in Fig 4. Our analysis demonstrates that MoM exhibits linear complexity, showcasing significant advantages over the Transformer model when handling long sequences. Specifically, MoM’s efficient memory update mechanisms allow it to process longer inputs with reduced computational overhead, positioning it as a more scalable solution for large-scale natural language processing tasks.
4.2.5 Length Extrapolation
We pretrained the models on the Slimpajama dataset with a 2K context length and conducted extrapolation experiments on various lengths using the Fineweb(penedo2024the) dataset. We extended the length to 32K to calculate perplexity (ppl). As shown in Fig 4, the Transformer model experienced a significant increase in ppl due to its poor extrapolation capability. Among the linear models, MoM achieved the best results.
[Uncaptioned image] Figure 3: Inference Efficiency of MoM. We demonstrate the inference time and GPU memory consumption required to generate 1K tokens at specific sequence lengths. [Uncaptioned image] Figure 4: Length Extrapolation. We extrapolated models trained on 2K sequences to a length of 32K for perplexity (ppl) evaluation.
4.2.6 Memory Analysis
Memory Load Balance Analysis. To evaluate whether each memory segment in MoM is effectively balanced during inference on downstream tasks, we analyzed the number of tokens routed to each layer using around 300k tokens from the ARC-easy benchmark. We visualized the results with auxiliary loss (following the formulation introduced in Switch Transformer fedus2022switch) in Fig 5 with heatmaps and we also visualized results with auxiliary loss in Fig 9. Due to the adoption of auxiliary loss, the memory segments in each layer are almost uniformly routed and activated.
Functions of Different Memories. We analyzed the memory routing within the model and identified the categories of tokens processed by each of the four memories. By examining an intermediate layer, we observed that, unlike traditional MoE in FFN layers where experts often lack clear specialization and focus on syntactic punctuation(jiang2023mistral), MoM displays specialization among its memories. This finding suggests potential for exploring an MoE architecture where each memory serves a specific role. These distinctions among different memories are presented in Table 5.
Table 5: Functions of Different Memories. By analyzing the types of tokens routed to the model’s intermediate layer, we observed a degree of specialization among the different memories.
Memory Index Type of Tokens Potential Function
1 Basic nouns/verbs/prepositions Simplify semantic information/causal logic
2 Proper nouns/scientific terms Specialized knowledge
3 Technical terms/adjectives Detailed memory
4 Interrogative words/incomplete nouns Fragmented/open information
Figure 5: Memory Load Balance Analysis. Token Routing Distribution Across Layers and Memories with Aux Loss.
4.2.7 MoM Scaling Up & Ablation Study
We examine the effect of scaling both the number of memory states and the number of top-$k$ activations in MoM. To ensure comparability, Fig. 6 reports results with a fixed activation ratio of 0.5. Increasing the number of memories from 1 to 8 consistently improves performance across both recall-intensive and commonsense benchmarks. These results indicate that enlarging the memory pool effectively mitigates interference and enhances capacity. More comprehensive results covering other activation ratios and activation settings are provided in Appendix H.
We further study the influence of auxiliary loss and shared memory in MoM, using a 380M-parameter model trained on 15B tokens. As shown in Table 6, auxiliary loss improves stability and performance when applied with a suitable weight. In addition, shared memory consistently benefits performance with global information. These results highlight the complementary roles of auxiliary loss and shared memory in stabilizing and enhancing MoM.
Figure 6: Scaling performance with increasing number of memories with a fixed activation ratio of 0.5.
Table 6: Ablation on memory count and shared memory, showing average accuracy across recall-intensive tasks.
Recall $\uparrow$ Common $\uparrow$
Aux Loss Scale
1e-2 27.59 42.10
5e-3 26.55 41.71
1e-3 28.16 41.97
0 27.23 41.58
Shared Memory
w/ shared memory 28.16 41.97
w/o shared memory 26.06 40.38
5 Conclusion
In this paper, we propose MoM, a novel architecture that enhances memory capacity and eliminates memory interference. By leveraging multiple independent memory states, MoM significantly improves performance on recall-intensive tasks while maintaining the efficiency advantages of linear models. Instead of simply discarding tokens as done in gating mechanisms, our memory separation paradigm provides a more effective way to preserve sequence information. Our experimental results demonstrate that MoM outperforms existing linear sequence modeling methods, particularly on tasks requiring strong recall, and achieves performance comparable to Transformer models. This makes MoM a promising approach for applications need strong efficiency and recall-intensive performance, paving the way for efficient sequence modeling.
6 Ethics Statement
This work does not involve human subjects, sensitive data, or high-risk applications. All experiments are conducted on publicly available datasets. We encourage responsible and ethical use of the proposed methods in line with community standards.
Appendix A Related Work
Linear Recurrent Models
Linear recurrent models, comprising linear attention, linear RNNs, state-space models (SSMs), have garnered significant research interests (qin2023scaling). The advancement of SSMs began with the pioneering work on S4 (gu2022efficientlymodelinglongsequences), which was later optimized through a diagonalized version (NEURIPS2022_9156b0f6). Despite their strong performance on the LRA benchmark, these models have faced challenges in language modeling mainly because they rely solely on data-independent processes. As research progressed, constant forgetting gates were introduced, helping to alleviate some interference issues by uniformly managing memory decay (sun2023retentive; gu2024mambalineartimesequencemodeling). The next breakthrough involved data-dependent forget gates. These allowed models to dynamically adjust memory updates based on the input data, significantly enhancing performance across various tasks (qin2024various; qin2024hgrn2; yang2023gated; zhang2024gated; yang2024gated; qin2024unlocking). Sequence parallelism techniques are well adapted on linear recurrent models (sun2024linear; sun2025lasp) for efficient long context training. There have also been recent advancements in scaling law and test-time regression optimization (shen2024scaling; sun2024learning; behrouz2024titans).
Building on these advancements, our MoM model incorporates data-dependent mechanisms that selectively update memory. By efficiently managing interference through tailored memory updates and leveraging increased memory capacity, MoM represents a further evolution, improving model expressiveness and performance.
Mixture-of-Experts
Mixture-of-Experts (MoE) is a technique designed to enhance the capacity of deep neural networks while maintaining computational efficiency (fedus2022switch; rajbhandari2020zero; lepikhin2020gshard; tang2023ms; zhu2024llama; qu2024llama). MoE achieves this by activating a subset of parameters, known as “experts”, for each input, which reduces the computational costs. Shazeer first integrated MoE into LSTM layers (shazeer2017outrageously). The Switch Transformer (fedus2022switch) refined this approach by simplifying the gating mechanism to select only one expert per input. Gshard (lepikhin2020gshard) further advanced this by using a top-2 expert routing strategy to improve performance. Recent MoE models, such as Deepseek-MoE (dai2024deepseekmoe), introduce shared experts to capture and consolidate common knowledge across different contexts, while designing fine-grained experts to increase combinatorial flexibility.
Appendix B Comparison between MoM and MoE
While our approach to implementing the Mixture-of-Memories (MoM) draws inspiration from the Mixture-of-Experts (MoE) framework, there are significant differences that distinguish our method from traditional MoE implementations.
  • Purpose: The MoE was introduced to scale up the number of parameters without significantly increasing computational resources. It address the limitations of dense models in scaling both parameters and computational demands through sparse activation. However, MoM is designed to expand the memory capacity of linear attention models while preserving their linear time complexity. By sparsely activating memories and using weighed summation to create a mixed memory, MoM effectively address the challenge of forgetting historical information in linear attention. Moreover, by separating the memory into distinct states, MoM reduces interference between different pieces of information.
  • Structure: In conventional MoE, each expert is a separate neural network within the feedforward network (FFN) layer such as Qwen-MoE (qwen_moe) and Linear-MoE sun2024linear. In contrast, in MoM, each memory is an RNN state with unique key-value projection weights to generate different key-value pairs. MoE operates during the channel mixing phase, where each token is processed independently by selected experts. On the other hand, MoM functions during the token mixing phase, where each memory processes different segments of the sequence, preserving inter-token relationships.
Appendix C Experiments Details
For the 380M models, we train on 15B tokens with a batch size of 0.5M tokens. The warmup tokens count is set to 0.25M. We set the hidden ratio of our model to 3 to keep the activated parameter count approximately the same. For the 1.3B models, we train on 100B tokens with a batch size of 2M tokens. The warmup tokens count is 1B. We employ AdamW optimizer (loshchilov2017fixing; sun2024co2) with learning rate of 3e-4 with cosine learning rate schedule (zhou2020pbsgd). The weight decay is set to 0.01 and gradient clipping is 1.0. Our experiments were conducted using 32 NVIDIA A800 GPUs. Training the 380M parameter model required approximately 10 hours, while the 1.3B parameter model took around 6 days.
Appendix D Commonsense Reasoning Tasks
As shown in Table 7, we report the language modeling perplexity and zero-shot performance of commonsense reasoning tasks following (zhang2024gated) which includes WikiText (merity2016pointer), LAMBADA (paperno2016lambada), ARC-easy, ARC-challenge (Clark2018ThinkYH), HellaSwag (zellers2019hellaswag), PiQA (Bisk2020) and WinoGrande (sakaguchi2019winogrande). The evaluation results are based on the lm-evaluation-harness (eval-harness).
Experimental results show that MoM outperforms other linear models and surpassed the Transformer model as well.
Table 7: Results on Common-Sense Reasoning Tasks. The performance of linear models and Transformer models is comparable; however, MoM consistently achieves the best average performance across all model sizes.
Scale Model
Wiki.
[…]
Lamb.
[…]
ARC-e
[…]
ARC-c
$\mathrm{acc_{n}}\uparrow$
Hella.
$\mathrm{acc_{n}}\uparrow$
Lamb.
$\mathrm{acc}\uparrow$
PIQA
$\mathrm{acc}\uparrow$
Wino.
$\mathrm{acc}\uparrow$
Avg.
380M Params 15B Tokens L=24, d=1024 Transformer++ 26.88 76.46 44.91 25.94 34.95 26.90 64.31 51.07 41.35
RetNet 31.07 87.11 44.49 23.04 33.86 23.93 63.49 52.33 40.19
HGRN2 27.90 77.40 45.24 23.63 35.61 24.74 65.45 54.06 41.46
GLA 28.78 79.95 44.53 22.27 34.84 24.94 63.93 51.38 40.32
GSA 28.17 82.50 45.50 24.23 35.00 24.02 64.85 50.43 40.67
Gated DeltaNet 26.47 58.59 46.04 23.55 35.18 27.01 66.05 50.83 41.44
MoM 25.86 55.41 44.65 24.74 36.54 27.93 66.16 51.78 41.97
1.3B Params 100B Tokens L=24, d=2048 Transformer++ 17.61 19.29 55.01 28.07 49.21 40.95 70.08 56.27 49.93
RetNet 18.18 21.97 57.49 26.88 48.09 37.75 69.37 53.28 48.81
HGRN2 17.32 15.65 58.33 28.07 51.93 42.31 71.33 52.01 50.66
GLA 17.61 19.66 55.18 27.56 48.89 40.03 69.86 53.91 49.24
GSA 16.69 16.02 58.33 28.33 50.98 42.03 72.25 53.43 50.89
Gated DeltaNet 17.14 18.80 56.82 27.39 49.77 39.94 71.76 51.78 49.58
MoM 16.64 14.83 55.35 27.99 50.95 43.43 71.27 56.83 50.97
Appendix E Training Loss Comparison
To further assess the learning efficiency of MoM, we compared the training loss curves of MoM with those of other baseline models. As depicted in Figure 7, MoM consistently maintains the lowest loss throughout the entire training phase. Even as training nears convergence, MoM continues to exhibit a clear advantage over other methods.
Figure 7: Training Loss. Loss curves for training 380M models on 15B tokens with a fixed random seed of 42.
Appendix F The Hybrid of MoM and Transformer
We delve deeper into the hybridization of MoM and Transformer layers by integrating 1 Transformer layer after every 7 MoM layers, resulting in only 3 Transformer layers across a total of 24 layers. The performance on commonsense reasoning and recall-intensive tasks is presented in the table 8. MoM-Hybrid demonstrates significantly improved results compared to Transformer models, despite using only 3 layers of global attention.
Table 8: Hybrid Model Performance. The hybrid model integrates 1 Transformer layer after every 7 MoM layers, resulting in only 3 Transformer layers across a total of 24 layers.
Model FDA SWDE SQUAD NQ TriviaQA Drop Avg.
Transformer++ 46.14 25.87 33.22 18.94 45.97 20.03 31.70
MoM 22.98 29.90 29.69 16.60 48.82 20.99 28.16
MoM Hybrid 58.13 44.05 35.71 20.18 48.10 20.60 37.80
Model
ARC-e
[…]
ARC-c
$\mathrm{acc_{n}}\uparrow$
Hella.
$\mathrm{acc_{n}}\uparrow$
Lamb.
$\mathrm{acc}\uparrow$
PIQA
$\mathrm{acc}\uparrow$
Wino.
$\mathrm{acc}\uparrow$
Avg.
Transformer++ 44.91 25.94 34.95 26.90 64.31 51.07 41.35
MoM 44.65 24.74 36.54 27.93 66.16 51.78 41.97
MoM Hybrid 46.55 24.49 36.45 28.86 65.51 52.41 42.38
Appendix G Fairness
To enhance memory capacity, MoM applies sparse activation to the key and value projections. Although these projections constitute a small portion of the overall model parameters, this inevitably increases the parameter count. Due to differing linear model structures, aligning both parameter count and memory capacity exactly is challenging. Thus, to ensure fairness, we conduct comparisons from two perspectives: equal activated parameter count and equal memory capacity.
G.1 Equal Activated Parameter Count
To ensure a fair comparison of parameters, we reduced the MLP hidden ratio to 2 and retrained the MoM model using the same training configurations as in Section 4.1. Both MoM and Gated Deltanet were set with 400M activated parameters. Although the smaller hidden ratio might impact the model’s commonsense knowledge, we tested on commonsense reasoning tasks and recall-intensive benchmarks. MoM consistently outperformed Gated Deltanet in both tests, further validating the effectiveness of the MoM approach. The results are presented in Table 9
Model Params
ARC-e
[…]
ARC-c
$\mathrm{acc_{n}}\uparrow$
Hella.
$\mathrm{acc_{n}}\uparrow$
Lamb.
$\mathrm{acc}\uparrow$
PIQA
$\mathrm{acc}\uparrow$
Wino.
$\mathrm{acc}\uparrow$
Avg.
Gated DeltaNet 400M 46.04 23.55 35.18 27.01 66.05 50.83 41.44
MoM 400M 47.10 23.72 35.43 26.88 64.64 51.22 41.50
Model Params FDA SWDE SQUAD NQ TriviaQA Drop Avg.
Gated DeltaNet 400M 20.53 23.24 28.55 14.98 44.91 16.48 24.78
MoM 400M 24.16 25.59 29.46 15.36 46.15 18.35 26.51
Table 9: Comparison with the Same Activated Parameters. MoM and Gated DeltaNet with 400M activated parameters are tested.
G.2 Equal Memory Capacity
To ensure a fair comparison of memory capacity, we also compared the single extended memory model with the MoM model. Notably, the single extended memory model has more parameters than the activated parameters in the MoM due to the extension of the $v$ dimension. MoM expands memory more elegantly and significantly outperforms in both recall and commonsense tasks. This comparison result is presented in Table 4.
Appendix H Detailed Scaling Results
To further examine the scalability of MoM from the perspective of memory capacity, we evaluate the effect of enlarging the memory pool beyond the main settings. Specifically, we compare two activation ratios, where the number of active memories accounts for either 0.5 or 0.25 of the total memory states. In all cases, an additional shared memory is included. Starting from a single memory as the baseline, we expand the number of memories to 2, 4, 8, and 16, and report the averaged results on both recall-intensive and commonsense benchmarks. The results, shown in Fig. 8, indicate that under a fixed activation ratio, increasing the memory size consistently improves performance on both categories of tasks.
Figure 8: Detailed scaling results. Performance shows a general improvement on both recall-intensive and commonsense tasks as the number of memories increases.
Appendix I Full Results
Model SQA MQA Sum FS Syn Code Zh-Avg En-Avg Avg.
RetNet 9.23 7.23 6.3 15.76 2.64 40.52 15.44 13.5 13.61
HGRN2 7.38 6.02 6.51 15.5 2.61 40.11 14.28 13.12 13.02
GSA 8.21 6.63 7.75 20.29 1.92 42.83 15.06 15.2 14.61
Gated DeltaNet 8.52 6.61 7.14 18 2.1 41.52 14.19 14.63 13.98
MoM 8.14 7.11 6.89 21.26 2.63 47.79 17.33 15.71 15.64
Table 10: Complete Results of LongBench. SQA: Single-doc QA, MQA: Multi-doc QA, Sum: Summarization, FS: Few-shot learning, Syn: Synthetic
Figure 9: Memory Load Balance Analysis. Token Routing Distribution Across Layers and Memories without Aux Loss.